OpenStudy (anonymous):

Please help!! 8. Solve these equations: a. sin(x + (π / 4)) – sin(x – (π / 4) = 1  b. sin x cos x = √3 / 4 c.tan^2 x–3tanx+2=0

OpenStudy (jack1):

b) i remember that \[\sin x \cos x = \frac 12 \sin (2x)\] does that help at all with this one?

OpenStudy (anonymous):

hmm I;m not sure. This is all very hard for me

OpenStudy (jack1):

\[\sin(x) \cos(x)= \frac 12 \sin(2x) = \frac {\sqrt 3}4\] \[\large \frac 12 \sin(2x) = \frac {\sqrt 3}4\] \[\large \sin(2x) = \frac {\sqrt 3}4 \times 2\] \[\large \sin(2x) = \frac {\sqrt 3}2\] \[\large (2x) = \sin^{-1} (\frac {\sqrt 3}2)\] \[\large 2x = \sin^{-1} (\frac {\sqrt 3}2)\] use calculator to work out from here, and tell me ur answer and I'll compare to mine? @washcaps ?

OpenStudy (anonymous):

pi/6?

OpenStudy (jack1):

perfect, that's what i got too... x = pi/6

OpenStudy (anonymous):

oh so that is the solution to b) then! what about the others?

OpenStudy (jack1):

i just tried a couple of different things for a)... none of them worked... @hartnn can u help explain this please, i'm lost

OpenStudy (jack1):

@ganeshie8 ur awesome at trig, do u have a sec?

hero (hero):

I can solve b. but it's kinda long.

OpenStudy (anonymous):

@Hero we solved be already

OpenStudy (jack1):

nah, i think we got b, a and c are still a mystery though

OpenStudy (jack1):

@iambatman ... u have a sec? help plz?

hero (hero):

c isn't hard. let tanx = y

hero (hero):

y^2 - 3y + 2 = 0

OpenStudy (jack1):

ahh, gotcha, then solve using quadratic equation? @Hero ? that's awesome!

hero (hero):

And then after you find the factors set equal to zero, then replace y with tan x again

hero (hero):

And finish solving

hero (hero):

set it equal to zero

OpenStudy (anonymous):

y=2,1

OpenStudy (anonymous):

@Hero

hero (hero):

Washcaps, what you do is this: (y - 2)(y - 1) = 0 y - 2 = 0 y - 1 = 0 tan(x) - 2 = 0 tan(x) - 1 = 0 tan(x) = 2 tan(x) = 1 then take the inverse tan of both sides to isolate x

OpenStudy (anonymous):

arctan x=2? = -2.18 +- 2pin

hero (hero):

For A use this formula: sin(A+B)=sin A cos B + cos A sin B sin(A-B)=sin A cos B - cos A sin B

OpenStudy (anonymous):

wait I still don't know c

hero (hero):

You don't know what x equals

hero (hero):

You have to take inverse tan of both sides to get: \(x = \tan^{-1}(2)\) \(x = \tan^{-1}(1)\)

OpenStudy (anonymous):

x=1.1, pi/4?

hero (hero):

Correct

OpenStudy (anonymous):

oh really? that's the solution? I thought it was more difficult

hero (hero):

Yes, that's the solution for c

OpenStudy (anonymous):

Do you know how to solve A)?

hero (hero):

For part a apply these formulas: sin(A+B)=sin A cos B + cos A sin B sin(A-B)=sin A cos B - cos A sin B

hero (hero):

I have an alternative solution for b that uses quadratic formula as well.

OpenStudy (anonymous):

I think I'm okay on b because I have an answer already (pi/6). could you walk me through a)? it is my last set of questions like this

hero (hero):

There's more than one solution for b though

OpenStudy (anonymous):

really?

hero (hero):

Yes, really.

OpenStudy (anonymous):

ok so I know pi/6 is one of them. what would be the other?

hero (hero):

I will use my steps and I will verify what the actual solutions are.

OpenStudy (anonymous):

Ok

hero (hero):

sin (x) cos (x) = √(3)/ 4 Multiply both sides by 4: 4sin(x)cos(x) = √3 Square both sides: (Remember \((ab)^c = a^cb^c\): (4sin(x)cos(x))^2 = √(3)^2 16sin^2(x)cos^2(x) = 3 16sin^2(x)(1 - sin^2(x)) = 3

hero (hero):

That's what we have so far. Now upon distribution: 16sin^2(x) - 16sin^4(x) = 3

hero (hero):

We will do a substitution and let sin^2(x) = y 16y - 16y^2 = 3

OpenStudy (anonymous):

so y=1/4, 3/4

hero (hero):

@washcaps...no

hero (hero):

After rearranging the terms we get: -16y^2 + 16y - 3 = 0 16y^2 - 16y + 3 = 0 Factoring this we get: 16y^2 -(12 + 4)y + 3 = 0 16y^2 - 12y - 4y + 3 = 0 3y(4y - 3) - 1(4y - 3) = 0 (4y - 3)3y - 1) = 0 4y - 3 = 0 3y - 1 = 0 *Important* ---> We now have to substitute the sin^2x back in for y: 4sin^2x - 3 = 0 3sin^2x - 1 = 0

hero (hero):

sin^2x = 3/4 sin^2x = 1/3 (sin(x))^2 = 3/4 (sin(x))^2 = 1/3 \(\sin(x) = \sqrt{\frac{3}{4}}\) \(\sin(x) = \sqrt{\frac{1}{3}}\)

hero (hero):

\(x = \sin^{-1}\left(\sqrt{\frac{3}{4}}\right)\) \(x = \sin^{-1}\left(\sqrt{\frac{1}{3}}\right)\)

OpenStudy (anonymous):

x=pi/3, 0.61

hero (hero):

We have to check negative results as well.

OpenStudy (anonymous):

how do we do that?

OpenStudy (ranga):

For a) use the identities: sin(A+B) = sin(A)cos(B) + cos(A)sin(B) sin(A-B) = sin(A)cos(B) - cos(A)sin(B) Therefore, sin(A+B) - sin(A-B) = 2cos(A)sin(B)

hero (hero):

@ranga, I already posted that earlier

OpenStudy (ranga):

sorry, didn't read the whole thing.

OpenStudy (anonymous):

@ranga I'm just not sure how to use the identity though

hero (hero):

x = A pi/4 = B

OpenStudy (anonymous):

@Hero can we finish b before we move to a please. I'm getting confused lol

hero (hero):

You take the inverse sine of the negative values \(\sin^{-1}\left(-\sqrt{\frac{3}{4}}\right)\) \(\sin^{-1}\left(-\sqrt{\frac{1}{3}}\right)\)

hero (hero):

The final thing is, now you have to plug the values you found for x back in to the original equation for b to make sure each value works.

OpenStudy (anonymous):

x=pi/6, 5pi/3, pi/3, and +- 0.61?

hero (hero):

Plug each of those values into b to make sure they work. You'll want to use exact values for this.

hero (hero):

You should only have four solutions, not five.

OpenStudy (anonymous):

5pi/3 doesn't work. it gives -sqrt 3/4

hero (hero):

Where'd you even get that from?

OpenStudy (anonymous):

what?

hero (hero):

5pi/3

OpenStudy (anonymous):

hmm I'm not sure lol. so the answers are pi/6, pi/3 and =- 0.61?

OpenStudy (anonymous):

0.6154 to be specific

hero (hero):

\(x = \sin^{-1}\left(\sqrt{\frac{3}{4}}\right)\) \(x = x = \sin^{-1}\left(\sqrt{\frac{1}{3}}\right)\) \(x = \sin^{-1}\left(-\sqrt{\frac{3}{4}}\right)\) \(x = x = \sin^{-1}\left(-\sqrt{\frac{1}{3}}\right)\)

hero (hero):

That's only four possible results

OpenStudy (anonymous):

sin^-1 (-sqrt 3/4) = 5pi/3

hero (hero):

Are you entering that into your calculator correctly

OpenStudy (anonymous):

yes. also I don't see where the pi/6 answer is

hero (hero):

I don't agree that you are entering that into your calculator correctly. Type how you are inputting it

hero (hero):

-sqrt(3/4) not -sqrt(3)/4

OpenStudy (anonymous):

that's what I typed. Could you just tell me the 4 solutions? I have to leave in 5 minutes and really need this done. You showed me the steps which i understand but arguing over the answers is taking too much time

hero (hero):

What calculator are you using?

OpenStudy (anonymous):

factoring calculator online

hero (hero):

give me the link to the site.

OpenStudy (anonymous):

http://www.freemathhelp.com/factoring-calculator.php please hurryy i have to leave

hero (hero):

How can you possibly calculate anything using that site? That's not a real calculator

OpenStudy (anonymous):

it seems to always work

hero (hero):

Explain how to use it

hero (hero):

It would be easier if you just tell me what you are inputting into it.

OpenStudy (anonymous):

I really don't have time. I just type in the equation and it gives me an answer. symbolab.com said the pi/6 is the only answer too. can you please just give me the 4 solutions. I will be back on in a couple hours

hero (hero):

Hang on

hero (hero):

As you can see, there are at least four values for x.

OpenStudy (anonymous):

I see that, but can't really read them proper;y. what are they?

hero (hero):

+/- 0.61 -pi/3 +pi/3 But the problem is, you have to check them because, from what I recall, only the two positive values work. But double check.

OpenStudy (anonymous):

jack1 and I arrived at pi/6 as an answer. read up top

hero (hero):

hang on. Let me show you something

OpenStudy (anonymous):

pi/6 is also true

hero (hero):

I know

hero (hero):

Those are the two correct answers then

OpenStudy (anonymous):

pi/3 and pi/6 are the solutions?